OpenStudy (anonymous):

solve for x (3x+1)/(x-5)=0 this is as far as i got from the original problem...

OpenStudy (lalaly):

3x+1=0 x=-1/3

OpenStudy (angela210793):

(3x+1)/(x-5)=0 only when (3x+1)=0 3x=-1 x=-1/3 also x should b different from 5

OpenStudy (anonymous):

Well in order for a fraction to be 0, the numerator must be 0, so just solve: 3x + 1 = 0 3x = -1 x = -1/3

OpenStudy (anonymous):

but what happened to the denominator?

OpenStudy (anonymous):

It doesn't matter what the denominator is, because you just want the numerator to be 0.

OpenStudy (anonymous):

multiply (x-5) for both sides so we can get rid of fractions

OpenStudy (anonymous):

ok, i see. thanks to all of you :)

OpenStudy (anonymous):

3x+1=x-5 2x=-4 x=-2

OpenStudy (lalaly):

asma nooo, il x-5 btroo7 ma3 il zero

OpenStudy (angela210793):

I didn't get wht Asma did O.o why did u equal tht?

OpenStudy (anonymous):

ahhhh sa777 nseeet 3ashan minzaman ma 3imilto 3 months without school lol!

OpenStudy (lalaly):

hehe its ok xP

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

sorry i did something wrong i just wanted to cancel the denominator and it is right bbut at the right side it have to be zero so sorry again

OpenStudy (angela210793):

ahhaaa....it's ok :D:D

OpenStudy (anonymous):

if P(x)/Q(X) = 0 <===> P(x) =0 , we solve only for the numerator

OpenStudy (anonymous):

=>3x+1=0 =>x=-1/3